# A 2.9-hour periodic radio transient with an optical counterpart

### v1

@@ 4 Conclusion
A binary system with an M-dwarf/neutron star is disfavoured by several arguments. If the radio emission is produced by a strong magnetic field, the hosted NS should be young, which is disfavoured by the high Galactic latitude. On the other hand, a scenario with a low-field NS in a low-mass binary (like the transitional pulsars or recycled pulsars), with the radio emission being dipolar spin-down powered, is disfavoured since the radio luminosity is several orders of magnitude larger than the limits on the spin-down.
@@
A binary system with an M-dwarf/white dwarf is a viable scenario. The radio emission in this case should be generated by a stellar wind flowing from the M-dwarf companion onto the magnetosphere of the white dwarf, where it is accelerated, producing radio emission. If this can be confirmed, these systems can be evolutionarily connected to AR Sco-type systems, intermediate polars, and polar WDs. A similar system might be the radio emitting polar candidate ILT J11015521 (de Ruiter et al., 2024).
@@
There is tentative evidence for a years periodicity possibly due to timing noise. If interpreted as an orbital period, it would place the source an order of magnitude too far apart to be interacting in a similar manner to AR Sco-type or ILT J11015521 systems. Furthermore, an M-dwarf/WD binary system with such a long orbital period would be challenging to form an evolutionary perspective. More data are needed to disentangle the nature of this long-timescale periodicity.

### v2

@@ 4 Conclusion
A binary system with an M-dwarf/neutron star is disfavoured by several arguments. If the radio emission is produced by a strong magnetic field, the hosted NS should be young, which is disfavoured by the high Galactic latitude. On the other hand, a scenario with a low-field NS in a low-mass binary (like the transitional pulsars or recycled pulsars), with the radio emission being dipolar spin-down powered, is disfavoured since the radio luminosity is several orders of magnitude larger than the limits on the spin-down power.
@@
A binary system with an M-dwarf/white dwarf is a viable scenario. The radio emission in this case should be generated by a stellar wind flowing from the M-dwarf companion onto the magnetosphere of the white dwarf, where it is accelerated, producing radio emission. If this can be confirmed, these systems can be evolutionarily connected to AR Sco-type systems, intermediate polars, and polar WDs. The 2.9 hr is most likely the spin and the orbit of a polar system, although other configurations cannot yet be completely excluded. A similar system might be the radio emitting polar candidate ILT J11015521 (de Ruiter et al., 2024).
@@
There is tentative evidence for a -year periodicity possibly due to timing noise. If interpreted as an orbital period, it would imply that the system components are separated by a distance greater by an order of magnitude than that which is seen in AR Sco-type or ILT J11015521 systems, making similar interactions unlikely. More data are needed to disentangle the nature of this long-timescale periodicity.